# Active learning emulators for nuclear two-body scattering in momentum space

### v1

@@ III.2.3 Emulation of total cross sections
Figure 5: Total npnp cross sections obtained with the G-ROM in panel (a) and LSPG-ROM in panel (b) (orange dots) based on the GT+ potential, which depends on nine LECs. The black lines depict the high-fidelity calculations. All partial-wave channels with $j\leqslant 6$ are included in the calculations. Both emulators use the same pool of candidate snapshots for training, comprising 500 random points selected via LHS. The insets show the mean relative error (23) of the emulator and the high-fidelity solution based on 1000 random samples in the potential’s parameter space (dashed lines) and the best-fit LEC values obtained in Ref. [Gezerlis:2014zia] (solid lines), for the requested error tolerances of $\alpha=10^{-4}$ and $10^{-7}$ (see the legends). See the main text for details.
Next, we benchmark our emulators by calculating the $np$ total cross section (9) as a function of the laboratory energy, using the GT+ potential. Figure 5 compares the emulator predictions (orange dots) and high-fidelity calculations (black lines). The results for the G-ROM are depicted in panel (a); for the LSPG-ROM in panel (b). All partial-wave channels with $j\leqslant 6$, coupled and uncoupled, are included in these calculations. The insets show the mean value of the relative error (23) of the emulators based on 1000 random samples in the potential’s nine-dimensional parameter space (dashed lines) and the best-fit LEC values obtained in Ref. [Gezerlis:2014zia] (solid lines). For emulator training, we generate a pool of 500 candidate snapshots, randomly selected via LHS within $\pm 40\%$ of the best-fit LEC values reported in Ref. [Gezerlis:2014zia]. Both ROMs have access to the same candidate snapshots during training.
@@ III.3 Proof of principle: Bayesian parameter estimation with emulator errors
Next, we explore the efficacy of our emulators incorporating error estimation for Bayesian calibration of chiral NN interactions. We perform these calculations here only as a proof-of-principle step toward future, more rigorous, applications in sophisticated Bayesian frameworks, such as those in Refs. [Melendez:2017phj, Melendez:2019izc, Wesolowski:2021cni, Svensson:2022kkj, Svensson:2023twt]. Our goal is to demonstrate that emulated errors for scattering observables can be estimated and accounted for in Bayesian calibration, and to assess the emulators’ computational efficiency.
To this end, we revisit the GT+ potential obtained in Ref. [Gezerlis:2014zia] and constrain its NN contact interactions using total cross-section data from the partial-wave analysis in Ref. [Stoks:1993tb]. However, let us emphasize again that our emulators are also applicable to more modern chiral NN interactions, including the LENPIC SMS potentials [Reinert:2017usi, Epelbaum:2022cyo]. Note that total cross sections are direct observables in contrast to the phase shifts.
@@ III.3.1 Statistical framework
we include the emulator error and assume the point-wise EFT truncation error model developed by the BUQEYE collaboration [Melendez:2020xcs]. We further assume that these uncertainties are independent and that the theory uncertainties dominate the uncertainty budget, i.e., $\delta y_{\text{exp}}(x)\ll\delta y_{\text{th}}(x;{\bf\it\theta})$. Future applications will more rigorously estimate the theory uncertainties, including correlated EFT truncation errors following the prescription in Ref. [Melendez:2019izc].
@@
The summation index $k$ runs over the coupled and uncoupled channels associated with a given total angular momentum $j$. To obtain Eq. (26), we assumed the conservative case in which the emulated on-shell component in $\mathbf{t}$ has maximum error, while the off-shell components (which do not contribute to $\sigma_{\text{tot}}$) have no error. As described in Sec. II.3, we also assume that $\norm{\mathbf{e}_{k}}\propto\norm{\mathbf{r}_{k}}$ and determine the factor of (approximate) proportionality for calibration as described in Ref. [Maldonado:2025ftg].
@@
We follow the point-wise truncation error model laid out in Appendix A in Ref. [Melendez:2019izc] to estimate the to-all-orders EFT truncation error,
@@
We assume that the reference scale $y_{\text{ref}}(x)$ is given by the EFT prediction at the best-fit values $\bf\it\theta$$^{*}$ obtained in Ref. [Gezerlis:2014zia], $y_{\text{ref}}(x)=y_{\text{th}}(x,{\bf\it\theta}^{*})$. At low energies, we can further assume that the EFT expansion parameter amounts to $Q=m_{\pi}/\Lambda_{b}$, where $m_{\pi}\approx 140\,\text{MeV}$ is the average pion mass and $\Lambda_{b}\approx 600\,\text{MeV}$ is the EFT breakdown scale [Melendez:2019izc]. Note that the $c_{n}(x)$ are not the LECs of the underlying interactions. In Eqs. (28), $\chi^{-2}(\nu_{0},\tau_{0}^{2})$ denotes the inverse-chi-squared distribution, with $\nu_{0}$ degrees of freedom and scale parameter $\tau_{0}^{2}$. As derived in Refs. [Melendez:2017phj, Melendez:2019izc], the EFT truncation error (28) conditional on $\bar{c}^{2}$ and $Q$ can then given by the normal distribution:
@@
indicating that the EFT truncation error is a fraction of $y_{\text{th}}^{2}(x,{\bf\it\theta}^{*})$. Here, we choose a realistic $\bar{c}=1$ [Millican:2025sdp, Wesolowski:2021cni, Melendez:2019izc].
@@
with the cumulative distribution function (CDF),
@@
The LEC prior distribution is taken to be an uncorrelated multivariate normal distribution with mean vector set to the best-fit values $\bf\it\theta$$^{*}$ obtained in Ref. [Gezerlis:2014zia]:
@@
with an arbitrarily chosen scaling factor $\kappa=0.1$ characterizing its width. The choices we have made for this proof-of-principle calculation can be straightforwardly adjusted using our source codes available on GitHub [BUQEYEsoftware].
@@ III.3.2 Inferred posterior distributions
Next, we perform Bayesian parameter estimation as outlined in the previous section using the emulator and the FOM solver separately, and then compare the results to assess the emulator’s efficacy. Since we found in Sec. III.2.3 that the G-ROM and LSPG-ROM behave very similarly in emulating total cross sections, we focus here on the G-ROM. The corresponding results for the LSPG-ROM are available in our GitHub repository [BUQEYEsoftware]. As in Sec. III.2.3, we train the G-ROM using the greedy algorithm initialized with two random snapshots and a pool of 500 random candidate snapshots to scan the nine-dimensional LEC space during the greedy iteration. We then let the greedy algorithm enrich the emulator basis until the requested error tolerance of $\alpha=10^{-2}$ is met, which must be done only once. Note that we have chosen a large tolerance $\alpha$ to explore the case in which the EFT truncation error (31) does not dominate the total theoretical uncertainty. This simulates a situation such as three-body scattering [Gnech:2025gsy, Gnech:2025lbg] where one can afford only relatively large emulator errors because otherwise the emulator training may be prohibitively slow.
@@
Figure 6: Subspace of the inferred posterior distribution (32) of the LECs in the GT+ potential optimized to the PWA93 total cross section [Stoks:1993tb]. The results based on the emulator (high-fidelity solver) are shown in red (blue), with the realistic $\bar{c}=1$ in Eq. (31). Note that the emulator error are purposely chosen large (by setting $\alpha=10^{-2}$) compared to the EFT truncation error to explore this special case. For brevity, we show here only a representative subspace of the nine-dimensional posterior distribution (32), consisting of the LECs $(C_{S},C_{T},C_{6},C_{7})$. The complete correlation plot is available in the accompanying GitHub repository [BUQEYEsoftware]. The blue lines and squares denote the best-fit values obtained in Ref. [Gezerlis:2014zia]. The titles of the marginal distributions (along the diagonal) and the dashed blue lines indicate the 95% credibility interval centered on the median. See the main text for details.
@@
For this specific case, Fig. 6 shows the results of our MC sampling of the LEC posterior distribution (32) based on the emulator (red lines) and high-fidelity solver (blue lines). All nine LECs of the GT+ potential are calibrated to total cross section data [Stoks:1993tb] at twelve laboratory energies, eleven evenly spaced in $E_{\text{lab}}=0.1\ldots 25.1\,\text{MeV}$ in increments of $2.5\,\text{MeV}$ (where the cross section changes rapidly with the energy) and an additional one at $60\,\text{MeV}$. All of these points are collected in $\mathcal{D}$.
@@
Specifically, Fig. 6 shows the correlation plot of a three-dimensional subspace of the inferred nine-dimensional LEC posterior (32). For brevity, we present only the representative subspace comprising $(C_{S},C_{T},C_{6},C_{7})$. The interested reader may find the complete correlation plot in the accompanying GitHub repository [BUQEYEsoftware]. The blue horizontal and vertical lines, as well as the associated squares, denote the best-fit values obtained in Ref. [Gezerlis:2014zia] using a different fit protocol, e.g., by fitting to scattering phase shifts, not the total cross section as done here. The titles of the marginal distributions (in the diagonal panels) report 95% credibility intervals (dashed blue lines) centered on each distribution’s median. For each random sample of the LEC posterior, we emulate the total cross section and check that the unitary violation of the $S_{\ell}$-matrix in all contributing partial-wave channels is $|1-\eta_{\ell}|\lesssim 10^{-2}$. As shown in Fig. 6, even with the purposely large chosen emulator error (compared to the EFT truncation error) the LEC inference is not significantly distorted.
@@ III.3.3 Computational efficiency of the emulators
Figure 7: Computational efficiency of the G-ROM (“G”) and LSPG-ROM (“LSPG”) in emulating the total cross section (9) based on partial waves with $j\leqslant 6$, shown as a function of the number of quadrature grid points used in the LS equation. All timings are carried out over four energies simultaneously: $E_{\text{lab}}=10,50,100,$ and $200\,\text{MeV}$. The computational efficiency is quantified by the speed-up factor (41) using our JAX implementation of both the FOM and the ROMs. Speed-up factors for the two ROMs and four different requested tolerances $\alpha$ for the emulated $\mathbf{t}$ are shown (see also the legend). The annotations report the results of least-squares fits to the data on a log-log scale to extract the scaling exponent. This metric is less implementation-dependent than the absolute speed-up values. Since the G-ROM and LSPG-ROM yield nearly identical basis sizes for a given tolerance, they exhibit very similar scaling behavior. For a typical grid size of $N=100$ (dashed vertical line), both achieve speed-up factors ranging from $S\approx 26$ to $S\approx 72$, depending on the tolerance. For the largest grid size of $N=120$, and the largest tolerance $\alpha=10^{-4}$, the speed-up increases further to $S\approx 92$. See the main text for details.
Figure 8: Basis sizes produced by the G-ROM emulator for each partial-wave channel at $E_{\text{lab}}=10\,\text{MeV}$ (left panel) and $E_{\text{lab}}=200\,\text{MeV}$ (right panel), shown for several requested tolerance values (see the legend) and a fixed number of momentum grid points $N=100$. At low energy, the required basis size decreases rapidly with increasing $j$, consistent with the suppression of high partial waves. At $E_{\text{lab}}=200\,\text{MeV}$, this decay is much weaker, and the high-$j$ channels require larger bases. As expected, stricter tolerances $\alpha$ lead to larger basis sizes for all channels. Coupled channels systematically require larger bases than uncoupled channels due to the increased dimensionality of the snapshots. The LSPG-ROM produces a similar pattern of basis sizes across channels, tolerances, and energies, closely mirroring the G-ROM results. See the main text for details.
Figure 9: The G-ROM and LSPG-ROM speed-up factor (41) follows a power-law dependence on the requested tolerance $\alpha$ for grid sizes $N=20,40,80$, and $120$, with larger grid sizes exhibiting slightly steeper exponents. Consequently, decreasing the tolerance by an order of magnitude reduces the speedup by roughly $26-35\%$. See Fig. 7 and the main text for details.
Having established the accuracy of the developed emulators, we now assess their computational efficiency. The efficiency of each ROM relative to the FOM is quantified by the speed-up factor,
@@
To obtain a clearer picture of the emulator performance and to provide an HPC-ready implementation suitable for Bayesian workflows, we implement both ROMs and the LS-based FOM solver in Python using JAX [JAX]. Leveraging an efficient offline-online decomposition together with just-in-time compilation and automatic vectorization offered by JAX, all potential matrix elements for all energies and partial-wave channels are precomputed and stored in the offline stage, as described in Sec. III.2. This procedure allows the FOM to evaluate the total cross section at many energies and channels in a single vectorized operation, providing an optimized baseline for comparing ROM speed-ups. Because the trained ROMs have different basis sizes across energy and partial-wave channels, we combine all individual $t$-matrix emulators into a unified array structure, enabling simultaneous evaluation in vectorized form. Finally, each ROM evaluation includes an on-the-fly estimate of the emulator error, which approximately doubles the per-evaluation run time; the reported speed-up factors therefore account for this additional cost.
Similar to Refs. [Garcia:2023slj, Maldonado:2025ftg], we emphasize that absolute speed-up factors are hardware- and implementation-dependent, and should therefore be interpreted with care. To mitigate this sensitivity, Fig. 7 examines the trend of the speed-up factor $S(N)$ as a function of the number $N$ of quadrature points. This scaling trend mainly reflects how much extra work each method requires when the problem becomes larger, which is a property of the method itself. As a result, the scaling curves are less affected by processor speed, memory layout, or compilation strategy, and the comparison becomes more general and easier to interpret.
@@
These basis-size trends play a central role in determining the overall computational cost of the ROM evaluations. As expected, decreasing the requested tolerance $\alpha$ systematically increases the number of snapshots for all channels, reflecting the tighter accuracy requirements imposed on the reduced representation. Additionally, because the reduced basis size depends strongly on the laboratory energy and the ROM evaluations are vectorized over channels and energies, the overall speed-up factor is typically limited by the coupled channels at the highest energy, which set the dominant computational bottleneck.
Figure 9 shows the emulator speed-up factor $S(\alpha)$ as a function of the requested ROM tolerance $\alpha$ for the momentum grid sizes $N=20,40,80$, and $120$. Across all $N$, the curves are nearly linear on a log-log scale, indicating a power-law dependence of the form $S(\alpha)\propto\alpha^{p}$, with $p$ ranging from 0.13(4) for $N=20$ up to 0.189(6) for $N=120$. This relationship means that tightening the tolerance by one order of magnitude reduces the speedup by a fixed multiplicative factor $10^{-p}$, corresponding to a fractional decrease of $1-10^{-p}$. The larger $N$ grids exhibit slightly larger exponents, and therefore a more pronounced decrease in speedup as the tolerance $\alpha$ becomes stricter. For example, for $N=20$, lowering the tolerance by a factor of 10 reduces the speedup by about $1-10^{-0.13}\approx 0.26$, or 26%. For $N=120$, the reduction is closer to 35%. Thus, while the grid must be sufficiently large to realize substantial speedups, the cost of tightening the tolerance also grows mildly with $N$. Even so, the emulator ultimately delivers its largest overall speedups at the largest grid sizes.

### v2

@@ III.2.3 Emulation of total cross sections
Figure 5: Total npnp cross sections obtained with the G-ROM in panel (a) and LSPG-ROM in panel (b) (orange dots) based on the GT+ potential, which depends on nine LECs. The black lines depict the high-fidelity calculations. All partial-wave channels with $j\leqslant 6$ are included in the calculations. Both emulators use the same pool of candidate snapshots for training, comprising 500 random points selected via LHS. The insets show the mean relative error (23) of the emulator and the high-fidelity solution based on 1000 random samples in the potential’s parameter space (dashed lines) and the best-fit LEC values obtained in Ref. [18] (solid lines), for the requested error tolerances of $\alpha=10^{-4}$ and $10^{-7}$ (see the legends).
Next, we benchmark our emulators by calculating the $np$ total cross section (9) as a function of the laboratory energy, using the GT+ potential. Figure 5 compares the emulator predictions (orange dots) and high-fidelity calculations (black lines). The results for the G-ROM are depicted in panel (a); for the LSPG-ROM in panel (b). All partial-wave channels with $j\leqslant 6$, coupled and uncoupled, are included in these calculations. The insets show the mean value of the relative error (23) of the emulators based on 1000 random samples in the potential’s nine-dimensional parameter space (dashed lines) and the best-fit LEC values obtained in Ref. [18] (solid lines). For emulator training, we generate a pool of 500 candidate snapshots, randomly selected via LHS within $\pm 40\%$ of the best-fit LEC values reported in Ref. [18]. Both ROMs have access to the same candidate snapshots during training.
@@ III.3 Proof of principle: Bayesian parameter estimation with emulator errors
Next, we explore the efficacy of our emulators incorporating error estimation for Bayesian calibration of chiral NN interactions. We perform these calculations here only as a proof-of-principle step toward future, more rigorous, applications in sophisticated Bayesian frameworks, such as those in Refs. [33, 32, 49, 45, 46]. Our goal is to demonstrate that emulated errors for scattering observables can be estimated and accounted for in Bayesian calibration, and to assess the emulators’ computational efficiency.
To this end, we revisit the GT+ potential obtained in Ref. [18] and constrain its NN contact interactions using total cross-section data from the partial-wave analysis in Ref. [44]. However, let us emphasize again that our emulators are also applicable to more modern chiral NN interactions, including the LENPIC SMS potentials [40, 13]. Note that total cross sections are direct observables in contrast to the phase shifts.
@@ III.3.1 Statistical framework
we include the emulator error and assume the point-wise EFT truncation error model developed by the BUQEYE collaboration [34]. We further assume that these uncertainties are independent and that the theory uncertainties dominate the uncertainty budget, i.e., $\delta y_{\text{exp}}(x)\ll\delta y_{\text{th}}(x;{\bf\it\theta})$. Future applications will more rigorously estimate the theory uncertainties, including correlated EFT truncation errors following the prescription in Ref. [32].
@@
The summation index $k$ runs over the coupled and uncoupled channels associated with a given total angular momentum $j$. To obtain Eq. (26), we assumed the conservative case in which the emulated on-shell component in $\mathbf{t}$ has maximum error, while the off-shell components (which do not contribute to $\sigma_{\text{tot}}$) have no error. As described in Sec. II.3, we also assume that $\norm{\mathbf{e}_{k}}\propto\norm{\mathbf{r}_{k}}$ and determine the factor of (approximate) proportionality for calibration as described in Ref. [28].
@@
We follow the point-wise truncation error model laid out in Appendix A in Ref. [32] to estimate the to-all-orders EFT truncation error,
@@
We assume that the reference scale $y_{\text{ref}}(x)$ is given by the EFT prediction at the best-fit values $\bf\it\theta$$^{*}$ obtained in Ref. [18], $y_{\text{ref}}(x)=y_{\text{th}}(x,{\bf\it\theta}^{*})$. At low energies, we can further assume that the EFT expansion parameter amounts to $Q=m_{\pi}/\Lambda_{b}$, where $m_{\pi}\approx 140\,\text{MeV}$ is the average pion mass and $\Lambda_{b}\approx 600\,\text{MeV}$ is the EFT breakdown scale [32]. Note that the $c_{n}(x)$ are not the LECs of the underlying interactions. In Eqs. (28), $\chi^{-2}(\nu_{0},\tau_{0}^{2})$ denotes the inverse-chi-squared distribution, with $\nu_{0}$ degrees of freedom and scale parameter $\tau_{0}^{2}$. As derived in Refs. [33, 32], the EFT truncation error (28) conditional on $\bar{c}^{2}$ and $Q$ can then given by the normal distribution:
@@
indicating that the EFT truncation error is a fraction of $y_{\text{th}}^{2}(x,{\bf\it\theta}^{*})$. Here, we choose a realistic $\bar{c}=1$ [36, 49, 32].
@@
with the cumulative distribution function,
@@
The LEC prior distribution is taken to be an uncorrelated multivariate normal distribution with mean vector set to the best-fit values $\bf\it\theta$$^{*}$ obtained in Ref. [18]:
@@
with an arbitrarily chosen scaling factor $\kappa=0.1$ characterizing its width. The choices we have made for this proof-of-principle calculation can be straightforwardly adjusted using our source codes available on GitHub [4].
@@ III.3.2 Inferred posterior distributions
Next, we perform Bayesian parameter estimation as outlined in the previous section using the emulator and the FOM solver separately, and then compare the results to assess the emulator’s efficacy. Since we found in Sec. III.2.3 that the G-ROM and LSPG-ROM behave very similarly in emulating total cross sections, we focus here on the G-ROM. The corresponding results for the LSPG-ROM are available in our GitHub repository [4]. As in Sec. III.2.3, we train the G-ROM using the greedy algorithm initialized with two random snapshots and a pool of 500 random candidate snapshots to scan the nine-dimensional LEC space during the greedy iteration. We then let the greedy algorithm enrich the emulator basis until the requested error tolerance of $\alpha=10^{-2}$ is met, which must be done only once. Note that we have chosen a large tolerance $\alpha$ to explore the case in which the EFT truncation error (31) does not dominate the total theoretical uncertainty. This simulates a situation such as three-body scattering [20, 21] where one can afford only relatively large emulator errors because otherwise the emulator training may be prohibitively slow.
@@
Figure 6: Subspace of the inferred posterior distribution (32) of the LECs in the GT+ potential optimized to the PWA93 total cross section [44]. The results based on the emulator (FOM solver) are shown in red (blue), with the realistic $\bar{c}=1$ in Eq. (31). For brevity, we show here only a representative subspace of the nine-dimensional posterior distribution (32), consisting of the LECs $(C_{S},C_{T},C_{6},C_{7})$. The blue lines and squares denote the best-fit values obtained in Ref. [18]. The titles of the marginal distributions (along the diagonal) and the dashed blue lines indicate the 95% credibility interval centered on the median.
@@
For this specific case, Fig. 6 shows the results of our MC sampling of the LEC posterior distribution (32) based on the emulator (red lines) and high-fidelity solver (blue lines). All nine LECs of the GT+ potential are calibrated to total cross section data [44] at twelve laboratory energies, eleven evenly spaced in $E_{\text{lab}}=0.1\ldots 25.1\,\text{MeV}$ in increments of $2.5\,\text{MeV}$ (where the cross section changes rapidly with the energy) and an additional one at $60\,\text{MeV}$. All of these points are collected in $\mathcal{D}$.
@@
Specifically, Fig. 6 shows the correlation plot of a three-dimensional subspace of the inferred nine-dimensional LEC posterior (32). For brevity, we present only the representative subspace comprising $(C_{S},C_{T},C_{6},C_{7})$. The interested reader may find the complete correlation plot in the accompanying GitHub repository [4]. The blue horizontal and vertical lines, as well as the associated squares, denote the best-fit values obtained in Ref. [18] using a different fit protocol, e.g., by fitting to scattering phase shifts, not the total cross section as done here. The titles of the marginal distributions (in the diagonal panels) report 95% credibility intervals (dashed blue lines) centered on each distribution’s median. For each random sample of the LEC posterior, we emulate the total cross section and check that the unitary violation of the $S_{\ell}$-matrix in all contributing partial-wave channels is $|1-\eta_{\ell}|\lesssim 10^{-2}$. As shown in Fig. 6, even with the purposely large chosen emulator error (compared to the EFT truncation error) the LEC inference is not significantly distorted.
@@ III.3.3 Computational efficiency of the emulators
Figure 7: Computational efficiency of the G-ROM (“G”) and LSPG-ROM (“LSPG”) in emulating the total cross section (9) based on partial waves with $j\leqslant 6$, shown as a function of the number of quadrature grid points used in the LS equation. All timings are carried out over four energies simultaneously: $E_{\text{lab}}=10,50,100,$ and $200\,\text{MeV}$. speedup factors for the two ROMs and four different requested tolerances $\alpha$ for the emulated $\mathbf{t}$ are shown (see also the legend). The annotations report the results of least-squares fits to the data on a log-log scale to extract the scaling exponent. For a typical grid size of $N=100$ (dashed vertical line), both achieve speedup factors ranging from $S\approx 26$ to $S\approx 72$, depending on the tolerance. For the largest grid size of $N=120$, and the largest tolerance $\alpha=10^{-4}$, the speedup increases further to $S\approx 92$.
Figure 8: Basis sizes produced by the G-ROM emulator for each partial-wave channel at $E_{\text{lab}}=10\,\text{MeV}$ (left panel) and $E_{\text{lab}}=200\,\text{MeV}$ (right panel), shown for several requested tolerance values (see the legend) and a fixed number of momentum grid points $N=100$. At low energy, the required basis size decreases rapidly with increasing $j$, consistent with the suppression of high partial waves. At $E_{\text{lab}}=200\,\text{MeV}$, this decay is much weaker, and the high-$j$ channels require larger bases. As expected, stricter tolerances $\alpha$ lead to larger basis sizes for all channels. The LSPG-ROM produces a similar pattern of basis sizes across channels, tolerances, and energies, closely mirroring the G-ROM results.
Figure 9: The G-ROM and LSPG-ROM speedup factor (41) follows a power-law dependence on the requested tolerance $\alpha$ for grid sizes $N=20,40,80$, and $120$, with larger grid sizes exhibiting slightly steeper exponents. Consequently, decreasing the tolerance by an order of magnitude reduces the speedup by roughly $26-35\%$. See Fig. 7 and the main text for details.
Having established the accuracy of the developed emulators, we now assess their computational efficiency. The efficiency of each ROM relative to the FOM is quantified by the speedup factor,
@@
To obtain a clearer picture of the emulator performance and to provide an HPC-ready implementation suitable for Bayesian workflows, we implement both ROMs and the LS-based FOM solver in Python using JAX [3]. Leveraging an efficient offline-online decomposition together with just-in-time compilation and automatic vectorization offered by JAX, all potential matrix elements for all energies and partial-wave channels are precomputed and stored in the offline stage, as described in Sec. III.2. This procedure allows the FOM to evaluate the total cross section at many energies and channels in a single vectorized operation, providing an optimized baseline for comparing ROM speedups. Because the trained ROMs have different basis sizes across energy and partial-wave channels, we combine all individual $t$-matrix emulators into a unified array structure, enabling simultaneous evaluation in vectorized form. Finally, each ROM evaluation includes an on-the-fly estimate of the emulator error, which approximately doubles the per-evaluation run time; the reported speedup factors therefore account for this additional cost.
Similar to Refs. [17, 28], we emphasize that absolute speedup factors are hardware- and implementation-dependent, and should therefore be interpreted with care. To mitigate this sensitivity, Fig. 7 examines the trend of the speedup factor $S(N)$ as a function of the number $N$ of quadrature points. This scaling trend mainly reflects how much extra work each method requires when the problem becomes larger, which is a property of the method itself. As a result, the scaling curves are less affected by processor speed, memory layout, or compilation strategy, and the comparison becomes more general and easier to interpret.
@@
These basis-size trends play a central role in determining the overall computational cost of the ROM evaluations. As expected, decreasing the requested tolerance $\alpha$ systematically increases the number of snapshots for all channels, reflecting the tighter accuracy requirements imposed on the reduced representation. Additionally, because the reduced basis size depends strongly on the laboratory energy and the ROM evaluations are vectorized over channels and energies, the overall speedup factor is typically limited by the coupled channels at the highest energy, which set the dominant computational bottleneck.
Figure 9 shows the emulator speedup factor $S(\alpha)$ as a function of the requested ROM tolerance $\alpha$ for the momentum grid sizes $N=20,40,80$, and $120$. Across all $N$, the curves are nearly linear on a log-log scale, indicating a power-law dependence of the form $S(\alpha)\propto\alpha^{p}$, with $p$ ranging from 0.13(4) for $N=20$ up to 0.189(6) for $N=120$. This relationship means that tightening the tolerance by one order of magnitude reduces the speedup by a fixed multiplicative factor $10^{-p}$, corresponding to a fractional decrease of $1-10^{-p}$. The larger $N$ grids exhibit slightly larger exponents, and therefore a more pronounced decrease in speedup as the tolerance $\alpha$ becomes stricter. For example, for $N=20$, lowering the tolerance by a factor of 10 reduces the speedup by about $1-10^{-0.13}\approx 0.26$, or 26%. For $N=120$, the reduction is closer to 35%. Thus, while the grid must be sufficiently large to realize substantial speedups, the cost of tightening the tolerance also grows mildly with $N$. Even so, the emulator ultimately delivers its largest overall speedups at the largest grid sizes.